# Social media polarization during conflict: Insights from an ideological stance dataset on Israel-Palestine Reddit comments

### v1

@@ Abstract
In politically sensitive scenarios like wars, social media serves as a platform for polarized discourse and expressions of strong ideological stances. While prior studies have explored ideological stance detection in general contexts, limited attention has been given to conflict-specific settings. This study addresses this gap by analyzing 9,969 Reddit comments related to the Israel-Palestine conflict, collected between October 2023 and August 2024. The comments were categorized into three stance classes: Pro-Israel, Pro-Palestine, and Neutral. Various approaches, including machine learning, pre-trained language models, neural networks, and prompt engineering strategies for open source large language models (LLMs), were employed to classify these stances. Performance was assessed using metrics such as accuracy, precision, recall, and F1-score. Among the tested methods, the Scoring and Reflective Re-read prompt in Mixtral 8x7B demonstrated the highest performance across all metrics. This study provides comparative insights into the effectiveness of different models for detecting ideological stances in highly polarized social media contexts. The dataset used in this research is publicly available for further exploration and validation.
Introduction
@@
Ideological stances on the part of one or both parties can increase the likelihood of conflicts between two nations. If the stances are sufficiently strong, they even render mutually acceptable transfers insufficient to prevent war [1]. And when a war has already broken out, these stances come to surface extensively on social media. [2] states that social media reinforce these existing positions through the fragmentation of the news media and facilitate the spread of misinformation. In the context of ongoing Israel-Palestine conflict, social media is a hub for the exchange of opposing political ideologies. And due to the historical nature of this conflict, this manifestation is often highly polarized from both sides.
@@
Numerous studies have been conducted in order to develop methodologies for identifying polarizing stances. Those studies have employed numerous machine learning algorithms, ranging from classical supervised algorithms to neural networks to pre-trained language models. Large Language Models (LLMs) have been used in sentiment detection of political texts [3]. A study [4] has also compared LLMs, neural networks, pre-trained language models and classical supervised algorithms in the context of ideological sentiment detection. Another study [5] compares LSTM and classical supervised algorithms in ideological stance detection. However, no studies have yet been conducted that use LLMs in ideological stance detection or compare LLMs to neural networks or pre-trained language models in this context. Moreover, no studies have also applied prompt engineering techniques in this context.
@@
To address this research gap, we present a comparative analysis of the performances of various open-access large language models (LLMs) incorporating multiple prompt engineering techniques, neural network architectures, and pre-trained language models in detecting ideological stance. We also propose the method that is identified as the best performing by the evaluation metrics. We specifically evaluated four open-access LLMs (Mixtral 8x7B, Mistral 7B, Gemma 7B and Falcon 7B), four neural network architectures (RNN, LSTM, GRU and BiLSTM), and seven pre-trained language models (BERT Cased, BERT Uncased, XLM-RoBERTa, DistilBERT Cased, DistilBERT Uncased, ELECTRA-Small and ELECTRA-Base). We utilized a manually annotated dataset of Reddit comments related to the Israel-Palestine conflict, spanning from October 2023 to August 2024. The models were evaluated based on accuracy, F1-score, recall, and precision on the test dataset. The final methodology was selected by systematically weighing these performance metrics. This comprehensive evaluation not only highlights the capabilities of various models in detecting ideological stance but also identifies the most suitable approach for this task. The contributions of this study are as follows:
@@
  • A meticulously manually annotated, publicly available dataset containing the ideological stances of 9,969 Reddit comments in the context of the Israel-Palestine conflict. [^1] The data is publicly available at https://github.com/jami78/Conflict-Bias-Eval
@@
  • A comprehensive comparative analysis of multiple models for ideological stance detection, particularly in politically sensitive and charged scenarios.
  • The establishment of benchmark models for this specific task, determined by evaluating and weighing various performance metrics.
@@
The remainder of the paper is organized as follows: Section Related Work provides a comprehensive review of the existing literature; Section Dataset outlines the details of the dataset; Section Methodology demonstrates the proposed methodology including the models that were used; Section Experimental Setup details the hardware setup and evaluation metrics used in our experiments; Section Performance Evaluation presents the performance scores of the models; and finally, Section Conclusion provides a conclusion with a summary of the findings and a few potential future directions.
Related Work
@@
Ideological stance detection in our context refers to the identification of the underlying ideological viewpoint that a text might have conveyed in response to a particular political scenario. Traditionally, this domain used to leverage classical machine learning techniques until the emergence of neural networks and pre-trained language models significantly changed the landscape. More recently, large language models (LLMs) have advanced this field along with several prompt engineering strategies as they can capture textual nuances that the previous models failed to do. This section reviews the progression of methodologies in ideological stance detection, organized into three categories: neural networks, pre-trained language models, and state-of-the-art LLMs. The previous studies of stance detection in the context of the Israel-Palestine conflict are also discussed.
Neural Networks
@@
Neural networks emerged as an improvement over the classical supervised model in the domain of ideological stance detection as it could capture patterns better. For example, [5] demonstrated that LSTM models outperform classical supervised models in ideological stance detection due to their ability to capture the temporal dependencies and context of tweets as observed in its performance on the Twitter dataset for 2019 Indian General Elections. According to [6], Bi-LSTM is even better than LSTM, RNN, and SVM with a motion-dependent debate model for this particular task. When an RNN model is enhanced with word embeddings (W2V), it outperforms traditional Logistic Regression that uses Bag of Words (BoW) and word embeddings (W2V) [7]. [8] proposes a BiGRU model with multiple embeddings that outperformed previous state-of-the-art models for target-dependent sentiment classification in news articles. The model’s success can be attributed to its ability to handle longer texts and multi-target sentences, offering a better solution for real-world sentiment analysis tasks in news media.
Pre-trained Language Models
@@
Pre-trained language models further improved performance over the neural networks by offering better handling of context. They have been used extensively over the years in ideological stance detection. According to [9], BERT demonstrates superior performance compared to classical supervised models, LSTM, and BiLSTM in this domain. However, [10] reported that a supervised SVM classifier with a motion-dependent debate model outperformed the BERT+MLP model in detecting ideological sentiment in a large UK parliamentary debate corpus known as ParlVote. Furthermore, ELECTRA, another pre-trained language model, has been shown to outperform BERT in the detection of ideological stance [9]. BERT, which utilizes transfer learning and logistic regression, has outperformed rule-based VADER and TextBlob in the detection of sentiment related to Sustainable Development Goals (SDGs) [11]. BERT-based models have also been shown to be more socially conservative compared to GPT models, effecting their classification performance downstream. As a result, these models might exhibit certain biases, particularly in hate speech detection, unless they are pre-trained on left-leaning or right-leaning corpora [12]. In a separate study, [13] concluded that language models such as BERT, RoBERTa, and DistilBERT continue to exhibit significant religious biases. RoBERTa-Base has also performed better than LSTM, VADER, XLM-R, and SVM in monolingual ideological sentiment classification [14]. However, according to [15], a RoBERTa-based model specifically fine-tuned for classification in the political domain, named POLITICS, outperformed RoBERTa-Base in terms of detecting ideological stance by leveraging discourse-based, structure-aware representations. Another multi-task learning model named CLoSE, introduced in [16], was shown to outperform both RoBERTa and BERT due to its capability of embedding sentence framing language and predicting ideological stance simultaneously. The study [17] showed that the TSM (Topic-Sentiment Matrix) model performed comparably to state-of-the-art distributional text representation models such as GloVe-d2v, achieving an accuracy of 65.04% compared to 64.30% with GloVe-d2v. DistilBERT has shown better performance than ELMo in the classification of protest news and sentiment analysis in cross-context settings [18].
Large Language Models (LLMs)
@@
Large Language Models (LLMs) are the latest innovation in stance detection which are capable of understanding contexts and textual nuances better than any previous model. For example- LLMs showed better classification performance than BERT, CNN, traditional supervised algorithms and neural networks in ideological sentiment classification [4]. Similarly, [3] evaluated the performance of several open-access large language models (LLMs) in detecting ideological sentiments which identified Falcon 40B as the best-performing model, followed by Mistral 7B and Falcon 7B. However, LLMs are not without challenges. LLMs have been shown to have a left-leaning tendency, which might affect their classification accuracy [19, 20]. Additionally, their high computational requirements might restrict their efficiency.
Stance Detection in the Israel-Palestine Conflict
@@
The Israel-Palestine war has instigated a few ideological stance detection studies, each with its own strengths and limitations. [21] investigated media stance in the coverage of the 2014 Israeli war on Gaza by Western news outlets. The study employed text mining techniques and supervised machine learning algorithms, such as SVM, to classify stances into binary categories of Pro-Palestine and Pro-Israel. The models showed high performance, achieving an accuracy and recall of approximately 90% in classifying these ideological stances. However, this study did not consider neutrality as a stance category, which limits its capacity to capture balanced stances in politically sensitive content. On the other hand, [22] incorporated neutrality as a stance category, broadening the range of perspectives considered. The study also achieved incredible model performance using XLNet. Nonetheless, it utilized a Twitter dataset, which may have limited the text’s ability to capture nuanced expressions due to the 280-character limit. Additionally, the study did not include state-of-the-art Large Language Models (LLMs) for its analysis.
@@
Building on this foundation, our work broadens the scope by integrating neural networks, pre-trained language models, and open source large language models (LLMs) with prompt engineering to analyze Reddit comments related to the Israel-Palestine conflict. Unlike traditional machine learning techniques, these advanced models enable a deeper understanding of contextual nuances, leading to the dynamic classification of ideological stance into three categories: Pro-Palestine, Pro-Israel, and Neutral.
Dataset
@@
The dataset used in this study is a filtered version of the ’Daily Public Opinion on Israel-Palestine War’ dataset [23], obtained from Kaggle. It contains comments from Reddit posts related to the ongoing Israel-Palestine war collected over a year from September 2023 to September 2024. This dataset has an advantage over previous datasets used for ideological stance detection because the context is set during a politically charged situation.
Data Collection and Filtering
The original dataset comprises 1,825,545 comments collected from various subreddits. In this study, we filtered the dataset using specific keywords to ensure contextual relevance for accurate classification. This filtering process resulted in a refined subset of 9,969 comments focused on terms such as “Hamas terrorist,” “Zionist terrorists,” “StandWithIsrael,” “FreePalestine,” and “Israel occupation,” among others. Subsequently, duplicate comments were removed to ensure that the dataset consisted of unique and unambiguous entries.
Data Annotation
@@
The labeling process utilized specific keywords for categorization:
  • Pro-Israel: Comments containing keywords that expressed support for Israel and opposition to Palestine or Hamas, such as “Hamas terrorist,” “StandWithIsrael,” “Hamas terror,” “SupportIsrael,” “IsraelStrong,” “IsraelForever,” “IsraelUnderAttack”.
  • Pro-Palestine: Comments containing keywords that expressed support for Palestine and opposition to Israel, such as “Zionist terrorists,” “Zionist terror,” “FreePalestine,” “boycottisrael,” “Palestine will be free,” “Palestinewillbefree”.
@@
To measure inter-annotator agreement, we utilized Fleiss’ Kappa [24]. The calculated score for the three annotators is 0.93, indicating a high level of agreement.
Dataset Statistics
As shown in Figure 1, “Hamas” and “Israel” emerge as the most frequently used keywords, surpassing others by a significant margin. The temporal distribution of comments (Figure 2) reveals that the number of comments is generally centered around 1,000 per month, with the highest volume observed in May 2024 and the lowest during the first and last two months of the dataset’s time span.
Fig 1: Top 5 keywords within the comments
Fig 2: Time Distribution of Comments
Data Processing

### v2

@@ Abstract
In politically sensitive scenarios such as wars, social media often becomes a platform for polarized discourse and expressions of strong ideological stances. While prior studies have explored ideological stance detection in general contexts, limited attention has been given to conflict-specific settings. This study addresses this gap by analyzing 9,969 Reddit comments related to the Israel–Palestine conflict collected between October 2023 and August 2024. The comments were manually annotated into three stance classes—Pro-Israel, Pro-Palestine, and Neutral—with high inter-annotator agreement (Fleiss’ Kappa = 0.93). We conduct a comprehensive comparative evaluation of neural networks, pre-trained language models, and large language models (LLMs), along with multiple prompt engineering strategies and an LLM majority voting ensemble. Model performance is assessed using accuracy, precision, recall, and F1-score. Among the tested approaches, the Scoring and Reflective Re-read prompting strategy applied to Mixtral 8x7B achieves the best overall performance across evaluation metrics. Beyond model performance, the findings provide empirical insights into how ideological polarization manifests in online discourse surrounding the Israel–Palestine conflict, offering a computational perspective on the distribution of Pro-Israel, Pro-Palestine, and Neutral narratives. The dataset used in this study is publicly available to support further research on ideological stance detection in politically sensitive contexts.
keywords:
Ideological Stance Detection, Machine Learning, Large Language Models, Israel-Palestine Conflict, Social Media Analysis
1 Introduction
@@
Ideological stances on the part of one or both parties can increase the likelihood of conflicts between two nations. If the stances are sufficiently strong, they even render mutually acceptable transfers insufficient to prevent war [jackson2007political]. And when a war has already broken out, these stances come to surface extensively on social media. [kubin2021role] states that social media reinforce these existing positions through the fragmentation of the news media and facilitate the spread of misinformation. In the context of recent Israel-Palestine conflict, social media is a hub for the exchange of opposing political ideologies. And due to the historical nature of this conflict, this manifestation is often highly polarized from both sides.
@@
Numerous studies have been conducted in order to develop methodologies for identifying polarizing stances. Those studies have employed numerous machine learning algorithms, ranging from classical supervised algorithms to neural networks to pre-trained language models. Large language models (LLMs) have been used in sentiment detection of political texts [kuila2024deciphering]. A study [kwon2024sentiment] has also compared LLMs, neural networks, pre-trained language models and classical supervised algorithms in the context of ideological sentiment detection. Another study [ansari2020analysis] compares LSTM and classical supervised algorithms in ideological stance detection. However, no studies have yet been conducted that use LLMs in ideological stance detection or compare LLMs to neural networks or pre-trained language models in this context. Moreover, no studies have also applied prompt engineering techniques in this context.
@@
To address this research gap, we conduct a comprehensive comparative analysis of multiple approaches for ideological stance detection, including open-access large language models (LLMs) with various prompt engineering techniques, neural network architectures, and pre-trained language models. We further identify the best-performing method based on multiple evaluation metrics. Specifically, we evaluate seven open-access LLMs (Mixtral 8x7B, Mistral 7B, Gemma 7B, Falcon 7B, Gemma 3-4B, Falcon 3-3B, and Qwen 3-4B), four neural network architectures (RNN, LSTM, GRU, and BiLSTM), and seven pre-trained language models (BERT Cased, BERT Uncased, XLM-RoBERTa, DistilBERT Cased, DistilBERT Uncased, ELECTRA-Small, and ELECTRA-Base). For LLM-based classification, we investigate several prompt engineering strategies, including Zero-shot prompting, One-shot prompting, Re2 prompting, Meta-prompting, Scoring and Reflective Re-read prompting, and Context Extraction prompting, and we also examine an LLM ensemble approach based on majority voting. We utilized a manually annotated dataset of Reddit comments related to the Israel-Palestine conflict, spanning from October 2023 to August 2024. The models were evaluated based on accuracy, F1-score, recall, and precision on the test dataset. The final methodology was selected by systematically weighing these performance metrics. This comprehensive evaluation not only highlights the capabilities of various models in detecting ideological stance but also identifies the most suitable approach for this task. The contributions of this study are as follows:
@@
  • A meticulously manually annotated and publicly available dataset containing the ideological stances of 9,969 Reddit comments in the context of the Israel-Palestine conflict, with high inter-annotator agreement (Fleiss’ Kappa = 0.93). [^1] The data is publicly available at https://github.com/jami78/Conflict-Bias-Eval
@@
  • A comprehensive comparative analysis of 18 models, including 7 pre-trained language models, 4 neural network architectures, and 7 large language model configurations, for ideological stance detection in politically sensitive and highly polarized scenarios.
  • The establishment of benchmark models and the development of novel prompt engineering strategies along with an LLM majority voting ensemble. Among the evaluated approaches, the proposed Scoring and Reflective Re-read prompting strategy achieves the best overall performance across evaluation metrics.
@@
The remainder of the paper is organized as follows: Section 2 provides a comprehensive review of the existing literature; Section 3 outlines the details of the dataset; Section 4 demonstrates the proposed methodology including the models that were used; Section 5 details the hardware setup and evaluation metrics used in our experiments; Section 6 presents the performance scores of the models; Section 7 analyzes the broader implications of this research and finally, Section 8 provides a conclusion with a summary of the findings and a few potential future directions.
2 Related Work
@@
Ideological stance detection in our context refers to the identification of the underlying ideological viewpoint that a text might have conveyed in response to a particular political scenario. Traditionally, this domain used to leverage classical machine learning techniques until the emergence of neural networks and pre-trained language models significantly changed the landscape [AlSarraj2018],[ansari2020analysis]. More recently, large language models (LLMs) have advanced this field with the help of several prompt engineering strategies due to their ability to capture textual nuances that the previous models failed to do [kuila2024deciphering], [mouthami2025]. This section reviews the progression of methodologies in ideological stance detection, organized into three categories: neural networks, pre-trained language models, and state-of-the-art LLMs. The previous studies of stance detection in the context of the Israel-Palestine conflict are also discussed.
2.1 Neural Networks
@@
Neural networks emerged as an improvement over the classical supervised model in the domain of ideological stance detection as it could capture patterns better. For example, [ansari2020analysis] demonstrated that LSTM models outperform classical supervised models in ideological stance detection due to their ability to capture the temporal dependencies and context of tweets as observed in its performance on the Twitter dataset for 2019 Indian General Elections. According to [khatua2020predicting], Bi-LSTM is even better than LSTM, RNN, and SVM with a motion-dependent debate model for this particular task. When an RNN model is enhanced with word embeddings (W2V), it outperforms traditional Logistic Regression that uses Bag of Words (BoW) and word embeddings (W2V) [iyyer2014political]. [hamborg2021news] proposes a BiGRU model with multiple embeddings that outperformed previous state-of-the-art models for target-dependent sentiment classification in news articles. The model’s success can be attributed to its ability to handle longer texts and multi-target sentences, offering a better solution for real-world sentiment analysis tasks in news media.
2.2 Pre-trained Language Models
@@
Pre-trained language models further improved performance over the neural networks by offering better handling of context. They have been used extensively over the years in ideological stance detection. According to [ozturk2024ideology], BERT demonstrates superior performance compared to classical supervised models, LSTM, and BiLSTM in this domain. However, [abercrombie2020parlvote] reported that a supervised SVM classifier with a motion-dependent debate model outperformed the BERT+MLP model in detecting ideological sentiment in a large UK parliamentary debate corpus known as ParlVote. Furthermore, ELECTRA, another pre-trained language model, has been shown to outperform BERT in the detection of ideological stance [ozturk2024ideology]. BERT, which utilizes transfer learning and logistic regression, has outperformed rule-based VADER and TextBlob in the detection of sentiment related to Sustainable Development Goals (SDGs) [rosenberg2023sentiment]. BERT-based models have also been shown to be more socially conservative compared to GPT models, effecting their classification performance downstream. As a result, these models might exhibit certain biases, particularly in hate speech detection, unless they are pre-trained on left-leaning or right-leaning corpora [feng2023from]. In a separate study, [abrar2025religious] concluded that language models such as BERT, RoBERTa, and DistilBERT continue to exhibit significant religious biases. RoBERTa-Base has also performed better than LSTM, VADER, XLM-R, and SVM in monolingual ideological sentiment classification [antypas2023negativity]. However, according to [ferracane2023integrated], a RoBERTa-based model specifically fine-tuned for classification in the political domain, named POLITICS, outperformed RoBERTa-Base in terms of detecting ideological stance by leveraging discourse-based, structure-aware representations. Another multi-task learning model named CLoSE, introduced in [kim2022close], was shown to outperform both RoBERTa and BERT due to its capability of embedding sentence framing language and predicting ideological stance simultaneously. The study [bhatia2018topic] showed that the TSM (Topic-Sentiment Matrix) model performed comparably to state-of-the-art distributional text representation models such as GloVe-d2v, achieving an accuracy of 65.04% compared to 64.30% with GloVe-d2v. DistilBERT has shown better performance than ELMo in the classification of protest news and sentiment analysis in cross-context settings [buyukoz2020analyzing].
2.3 Large Language Models (LLMs)
@@
Large language models (LLMs) are the latest innovation in stance detection which are capable of understanding contexts and textual nuances better than any previous model. For example- LLMs showed better classification performance than BERT, CNN, traditional supervised algorithms and neural networks in ideological sentiment classification [kwon2024sentiment]. Similarly, [kuila2024deciphering] evaluated the performance of several open-access large language models (LLMs) in detecting ideological sentiments which identified Falcon 40B as the best-performing model, followed by Mistral 7B and Falcon 7B. However, LLMs are not without challenges. LLMs have been shown to have a left-leaning tendency, which might affect their classification accuracy [hernandes2024llms, pit2023whose]. Additionally, their high computational requirements might restrict their efficiency.
2.4 Stance Detection in the Israel-Palestine Conflict
@@
The Israel-Palestine war has instigated a few ideological stance detection studies, each with its own strengths and limitations. [AlSarraj2018] investigated media stance in the coverage of the 2014 Israeli war on Gaza by Western news outlets. The study employed text mining techniques and supervised machine learning algorithms, such as SVM, to classify stances into binary categories of Pro-Palestine and Pro-Israel. The models showed high performance, achieving an accuracy and recall of approximately 90% in classifying these ideological stances. However, this study did not consider neutrality as a stance category, which limits its capacity to capture balanced stances in politically sensitive content. On the other hand, [imtiaz2022takingsidespublicopinion] incorporated neutrality as a stance category, broadening the range of perspectives considered. The study also achieved incredible model performance using XLNet. Nonetheless, it utilized a Twitter dataset, which may have limited the text’s ability to capture nuanced expressions due to the 280-character limit. Additionally, the study did not include state-of-the-art large language models (LLMs) for its analysis.
@@
Building on this foundation, our work broadens the scope by integrating neural networks, pre-trained language models, and open source large language models (LLMs) with prompt engineering to analyze Reddit comments related to the Israel-Palestine conflict. Unlike traditional machine learning techniques, these advanced models enable a deeper understanding of contextual nuances, leading to the dynamic classification of ideological stance into three categories: Pro-Palestine, Pro-Israel, and Neutral. Although [kuila2024deciphering] applied several existing prompting techniques for political sentiment classification, they did not introduce any novel methods for superior stance detection performance. We address this by evaluating multiple novel prompting strategies and proposing a new technique that improves stance detection on a manually annotated Reddit dataset of the Israel-Palestine conflict, which captures longer and more nuanced discourse than typical Twitter [imtiaz2022takingsidespublicopinion] and non-conflict datasets.
3 Dataset
@@
The dataset used in this study is a filtered version of the ’Daily Public Opinion on Israel-Palestine War’ dataset [asaniczka2024daily], obtained from Kaggle. It contains comments from Reddit posts related to the ongoing Israel-Palestine war collected over a year from September 2023 to September 2024. This dataset has an advantage over previous datasets used for ideological stance detection because the context is set during a politically charged situation.
3.1 Dataset Comparison
Table 1 illustrates the difference between the previous datasets and ours. Our dataset has an advantage over them due to the context being set during a politically charged situation. It is also manually annotated to a greater degree than the previous ones.
Table 1: Comparison of related datasets
Dataset Platform Conflict/Topic Year Size Stance Classes Limitations
Imtiaz et al. (2022) [imtiaz2022takingsidespublicopinion] Twitter Israel-Palestine Conflict 2021 493,877 tweets Pro-Palestine, Pro-Israel, Neutral Short-form tweets; limited manual annotation
Al-Sarraj et al. (2018) [AlSarraj2018] News websites Israel-Palestine Conflict 2014 799 articles Pro-Palestine, Pro-Israel No neutral class; limited time span
Ansari et al. (2020) [ansari2020analysis] Twitter Indian General Elections 2019 3,896 tweets Multi-party labels Not conflict-focused
Hernandez et al. (2022) [hernandes2024llms] News articles U.S. political bias 5,877 articles Multi-level ideology Outlet-based labels
Bhatia et al. (2018) [bhatia2018topic] Parliamentary debates U.S. ideology 2005 1,586 docs Republican, Democrat Two-party only
Ours Reddit Israel-Palestine Conflict 2023–2024 9,969 comments Pro-Palestine, Pro-Israel, Neutral Conflict-specific
3.2 Data Collection and Filtering
The original dataset comprises 1,825,545 comments collected from various subreddits. In this study, we used keyword filtering, a widely used NLP technique for summarization and retrieval, in order to identify contextually relevant comments as summarized by [chen2024isamasredpublicdatasettracking]. This filtering process resulted in a refined subset of 9,969 comments focused on terms such as “Hamas terrorist”, “Zionist terrorists”, “Stand With Israel”, “Free Palestine”, “October 7”, “From the river to the sea”, “Jewish state”, “Hostage”, and “Israel occupation”. This sample size strikes a balance between representativeness (ensuring that the comments capture complete and diverse expressions) and feasibility (enabling high-quality manual annotation and systematic LLM evaluation). Subsequently, duplicate comments were removed to ensure that the dataset consisted of unique and unambiguous entries.
3.3 Data Annotation
@@
The process utilized an initial filtering process using specific keywords before annotation:
  • Pro-Israel: Comments containing keywords that expressed support for Israel and opposition to Palestine or Hamas, such as “Hamas terrorist”, “Stand With Israel”, “Hamas terror”, “Support Israel”, “Israel Strong”, “Israel Forever”, “Israel Under Attack”.
  • Pro-Palestine: Comments containing keywords that expressed support for Palestine and opposition to Israel, such as “Zionist terrorists”, “Zionist terror”, ”Free Palestine”, “Boycott Israel”, “Palestine will be free”.
@@
To measure inter-annotator agreement, we utilized Fleiss’ Kappa [fleiss1971measuring]. The calculated score for the three annotators is 0.93, indicating a high level of agreement.
3.4 Dataset Statistics
Table 2 shows dataset statistics. There are a total of 5589 users with 1.78 average comment each. The entire dataset contains a total of 28,116 unique words and 12 unique subreddits. Among the subreddits, r/IsraelPalestine, r/worldnews, and r/Palestine are the most frequent as shown in figure 1.
Table 2: Dataset Overview
Metric Value
Total comments 9,969
Unique users 5,589
Comments per user (mean) 1.78
Total words (tokens) 1,149,661
Unique words (vocabulary size) 28,116
Unique subreddits 12
Figure 1: Top 7 Subreddits
According to Figure 2, “hamas”, and “israel”, emerge as the most frequently used keywords, surpassing others by a significant margin. Interestingly, these two words are also the most frequently used keywords in the texts corresponding to all three stances (Figure  3). However, the word “terrorist” appears significantly more in Pro-Israel stances than the other two. The word “gaza” also appears regularly in all three stances.
The temporal distribution of comments (Figure 4) reveals that the number of comments is generally centered around 1,000 per month, with the highest volume observed in May 2024 and the lowest during the first and last two months of the dataset’s time span. The original, unfiltered dataset does not show the exact same trend since November 2023 contains the highest number of comments according to figure 5. In terms of stance proportion per month, Pro-Palestine had highest number in September 2023 which dipped below the Pro-Israel level in October 2023. After that, the proportion remained consistent (Figure 6).
Figure 2: Top 10 Keywords
Figure 3: Keywords Per Stance
Figure 4: Time Distribution of Comments
Figure 5: Monthly Comment Volume of the Original Dataset
Figure 6: Stance Proportions Over Time
3.5 Data Processing